OpenStudy (anonymous):

Find the first five terms; a1=-5, an = 4a n-1 +10

OpenStudy (bibby):

do you see the relationship? write out the first 3 terms a1 = -5 a2 = a1 + 10 = -5 + 10 = 5 a3 = a2 + 10 = 5 + 10 = 15

OpenStudy (superdavesuper):

@bibby isnt the eqn: a_n = 4a_n-1 +10 ?

OpenStudy (bibby):

I can't read. whoops. a1 = -5 a2 = 4(a1)+10 = 4(-5) + 10 = -20 + 10 = -10 Try the rest

OpenStudy (superdavesuper):

yup :) so a3=4(a2)+10=4(-10)+10=-30

OpenStudy (anonymous):

ok so i got a4=4(a3)+10=4 thats is what i have so far..

OpenStudy (anonymous):

not sure what next

OpenStudy (superdavesuper):

hmmm....try to calculate a4 again plz? math seems off....

OpenStudy (superdavesuper):

remember a3=-30 :)

OpenStudy (anonymous):

a4=-110?

OpenStudy (superdavesuper):

Correct @lashae1995 :) i bet u can do a5 as well lol

OpenStudy (anonymous):

a5=-430?

OpenStudy (bibby):

Yep. -440+10.

OpenStudy (superdavesuper):

Good work! :)